Example of the L^{A}TEX-Package ieeepes for IEEE PES Transactions

V. Kuhlmann

Dept of Electrical and Electronic Engineering Christchurch, New Zealand Second Author affiliation

Abstract—Put the text of your abstract here.

I. Section

Text for the first section.

A. Subsection

Text for the first sub-section.

B. Subsection

Text for the second sub-section.

1) Subsubsection: Text for the first sub-sub-section.

2) Subsubsection: Text for the second sub-sub-section.

C. Subsection

Text for the third sub-section.

II. Section

Text for the second section.

figure matter

Fig. 1. This is the caption for figure #1. Make sure it goes after the figure!

Table I This is the caption for table #1. Make sure it goes before the table!

table matter

III. Figures and Tables

Text for the third section. This section has figures and tables in it.

And more text in this section. Figure and table references: Fig. 1, Table I. Use these

at the beginning and within a sentence.

Much better is to use the Figure and Table environments, which will take care of placing the caption correctly for you. See Fig. 2 and Table II.

Using the Figure environment:

The figure matter goes here.

Fig. 2. Caption for figure, Figure environment.

text Using the Table environment: And more text in this section. And more text in this

section. And more text in this section. And more text in this section. And more text in this section.

IV. Equations

Referencing equations: (1) for whatever. Equation (1) at the beginning of a sentence.

equation  (1)

text

V. References

VI. Test of Biographies

text

Table II Caption for table, Table environment.

The table matter goes here.

Author 1 text

text

Author 2 text

text

Author 3 text text

DRAFT  1999/6/1  2  1999/6/1  DRAFT

text 1 2 Discussion of PAPER NUMBER, “PAPER TITLE”, by AUTHOR NAMES. Discussion by DISCUSSER NAME, AFFILIATION INCL ADDRESS.

Discussion

DISCUSSER NAME (SHORT AFFILIATION) text

Closure

AUTHOR NAME text